Tuesday 04/29/14

9:30a – 10:00p

Computer Cajal troubleshooting

  • computer behaving very sluggish today, denying connections
  • restarted Cajal.
  • also removed all the memory DIMMs and reinstalled them. Cajal now recognizes 48 GB of RAM again (instead of just 24 GB).

Project 2

  • moved system back down-stairs to STORM4

STORM2 storm-control

  • added project 2 branch and jeff branch back on STORM2. These should run the new Dave and my modified daxwriters without issue.
  • To do:
    • test project 2 branch. Branch this into a chromatin branch or something.
    • update the project 2 daxwriters to do the adaptive scaling (normalized and rescaled to 2**15)
    • push updated and tested branch back to my git storm-control
  • Steve not running on project 2 branch. Need to update xml.

Chromatin Project

Cell staining

  • 2 10 min hot washes for new samples (3 2-clr and 2 1-clr, see yesterday’s post).

Data Analysis

  • ChromatinCropper Analyzed more C03 spots up through image 21. (still more data to process if desired)
  • Running bead fitting for C04 and C05 data
  • finishing running C02 analysis
  • Analyzing C04 with ChromatinCropper. some slides still not imaged.
  • Analyzed C05 with ChromatinCropper up to image 34. (50+ cells)
  • Analyzed C06 with ChromatinCropper up to image 25 (50 images)
  • C07 not finished running data processing with DaoSTORM. Now running on Cajal AND tuck (oops!)
  • C08 not finished running data processing with DaoSTORM. Now running on Cajal

More bead analysis

  • Analyzing bead data for E01, E02 and E03

Data copying

  • E09 and F07 data still just on AlistairTemp2. Now copying over slowly to RAID2 (which is also undergoing much data analysis, hence the very slow copy speed).

Bogdan black data

  • D11 first attempt imaging: high background, over-diluted beads.
  • 150K frames, still little 405 added.
  • try 1 hr 60C wash.
  • try cutting concentration 2 or 5 fold (was ~5,000 ng of probe in 20 uL)

STORM2 imaging

  • issues with microscope control software — project 2 branch not opening steve, storm2 branch skipping tiles.
  • fixed additional issue with field of view in hal transposed. Fixed on both project 2 branch and storm2 branch. did not push back to release (wait until Hazen troubleshoots other issues).
  • Imaging G08-P3 G09-P1 sample
  • in “fresh” MEA (MEA stock going bad / turning liquid).
  • very high background in P3 / 750 channel. should use less probe? wash more?
  • performed coarse alignment with bead brightness on camera and spot direction in epi mode to attempt to get the beam more centered on the sample. Traditional centering the beam spot on the camera leads to clearly off-center illumination, brightest in the upper right and dimmest in the lower left of the field of view.
This entry was posted in Summaries and tagged , , . Bookmark the permalink.